Three consecutive odd integers are such that the square of the third is 264 more than the square of the second. Find the three integers.: Three consecutive odd integers are such that the square of the third is 264 more than the square of the second. Find the three integers.
Answer by mangopeeler07(442)

consecutive odd integers= x, x+2, x+4
(x+4)^2=(x+2)^2+264

Simplify
x^2+8x+16=x^2+4x+4+264

Combine like terms on each side
x^2+8x+16=x^2+4x+268

Subtract 4x from both sides
x^2+4x+16=x^2+268

Subtract 268 from both sides
x^2+4x-252=x^2

Subtract x^2 from both sides
4x-252=0

Add 252 to both sides
4x=252

Divide both sides by 4
x=63

consecutive odd integers= x, x+2, x+4
----------------------------------------------------------------------------
Answer: 63, 65, and 67